OpenStudy (anonymous):

The area of the trapezoid below is 72 square inches. What is the length of the altitude DE?

OpenStudy (campbell_st):

do you know the lengths of the parallel sides...?

OpenStudy (acxbox22):

\(\Huge\color{purple}{\bigstar}\Huge\color{purple}{\bigstar}\Huge\color{purple}{\bigstar}\Large\color{gold}{WELCOME~TO~OPENSTUDY}\Huge\color{purple}{\bigstar}\Huge\color{purple}{\bigstar}\Huge\color{purple}{\bigstar}\)

OpenStudy (acxbox22):

where is the trapezoid below?

OpenStudy (campbell_st):

so using the formula \[72 = \frac{h}{2}\times (a + b)\] h = height and a and b are the lengths of the parallel sides which can be written as \[144 = h \times (a + b)\] then \[\frac{144}{(a + b)} = h\]

OpenStudy (anonymous):

The top of it is 14 and the bottom of it is 22.

OpenStudy (acxbox22):

A=(b1+b2)h/2 A=(22+14)h/2 72=36h/2 144=36h --------- 36 height/altitude=4 units

OpenStudy (campbell_st):

so substitute the values for a and b into the modified form of the area fromula to find the height